# Comprehensive Review and Empirical Evaluation of Causal Discovery Algorithms for Numerical Data

### v1

@@ 2.4 Relevant Survey on Causal Analysis
Figure 4: Question-Data-Method-Evaluation Template
Drawing upon the analysis mentioned above, we have amalgamated and organized the principal research directions, which are visually represented in Figure 4. By employing the Question-Data-Method-Evaluation (QDME) template, the review aims to provide a clear and structured overview of the diverse areas and subtopics within the field of causal discovery research, enhancing the organization and coherence of the work.
Based on the aforementioned papers, while the algorithms and evaluations of causal discovery have become relatively comprehensive, several unresolved challenges persist, highlighting gaps in the existing body of knowledge. First, many articles employ outdated taxonomies and lack updates on the latest algorithms. Second, most surveys emphasize theoretical analyses, often neglecting the systematic experiments necessary for quantitative assessment. Third, although some articles have conducted experimental comparisons of algorithms, these studies primarily consider causal structures as their experimental factors, overlooking the characteristics of the data. Consequently, there is currently an absence of evaluation systems that consider the user’s perspective.
@@ 3.1 Method Design
The graph illustrates an escalating annual trend in research endeavours about causal discovery. This discernible growth can be attributed partly to the robust advancement of AI technologies in contemporary times, which has laid the groundwork for algorithmic developments facilitated by enhanced data processing capabilities.
@@ 3.2 Data Collection
‘Causal Discovery” + [Survey |||| Overview |||| Review] 980 15 14
@@
Table 2: Datasets Collection Result
@@
Table 3: Metrics Collection Result
@@
3.3 Survey Analysis
@@ 4 Causal Discovery Algorithms
Methods, align=center [Information Theoretic-Based Approach - oCSE (Sun et al., 2015) ] [Causal Network-Based
Approach, align=center, s sep=2pt, l=0.9cm [Constraint-Based, [Peter-Clark (PC) (Kalisch and Bühlman, 2007)] [CD-NOD (Huang et al., 2020)] [PCMCI (Runge et al., 2019)] [Fast Causal Inference (FCI) (Entner and Hoyer, 2010b, a)] ] [Score-Based, [Greedy Equivalence Search (GES) (Chickering, 2002)] [ES (Yuan and Malone, 2013)] [GRaSP (Lam et al., 2022)] [DYNOTEARS (Pamfil et al., 2020)] ] ] ] [State Space
@@
Modelling-Based
Methods, align=center [Structural Equation Modeling (SEM) (Shimizu et al., 2006)] [LiNGAM-Based, [ICA-LiNGAM (Zhang et al., 2021a)] [Direct LiNGAM (Shimizu et al., 2011)] [VARLiNGAM (Hyvärinen et al., 2010)] [RCD (Maeda and Shimizu, 2020)] [CAM-UV (Maeda and Shimizu, 2021)] ] [Additive Noise Models (ANM) (Hoyer et al., 2008), [TiMINo (Peters et al., 2013)] ] [PNL (Zhang and Hyvarinen, 2012)] [DAGs with NO TEARS (Zheng et al., 2018)] [GOLEM (Ng et al., 2020)] ] [Deep Learning-Based
Methods, align=center [Causal Generative Neural Networks (CGNN) (Goudet et al., 2018)] [DAG-GNN (Zhang et al., 2021a)] [Temporal Causal Discovery Framework (TCDF) (Nauta et al., 2019a)] [Amortized Causal Discovery (ACD) (Löwe et al., 2022)] [Ordering-Based Causal Discovery with Reinforcement Learning (Wang et al., 2021)] [GraNDAG (Lachapelle et al., 2019)] ] [Hybrid Methods, align=center [ARMA-LiNGAM (Kawahara et al., 2011)] [Scalable Causation Discovery Algorithm (SCDA) (Raghu et al., 2018)] [PSDR-TE (Mao and Shang, 2017)] [Information Geometric Causal Inference (IGCI) (Janzing et al., 2012)] [Neural Granger Causality (NGC) (Tank et al., 2021b)] ] ]
Figure 6: Taxonomy for Causal Discovery (The underlined algorithm is only applicable to Time-Series)
@@ 4.1 Granger Based Method
Our attention is directed towards several primary algorithms based on GC. Arize et al. (1993) put forth the Multivariate Granger (MVGC) analysis method to overcome the limitations of Pairwise Granger (PWGC), which can only deal with bivariate data. However, ensuring linearity in real-world datasets can present a significant challenge. To address this issue, Chen et al. (2004) introduced an approach known as Extended Granger Causality (EGC), specifically designed to handle nonlinear data. An alternative model catering to nonlinear time series is the Kernel Granger (KGC) method (Marinazzo et al., 2008; Liao et al., 2009; Marinazzo et al., 2011), showcasing notable attributes such as high accuracy and flexibility. When confronted with continuous time series data, Hu et al. (2014) proposed the Copula Granger method, which is capable of uncovering nonlinear and higher-order causal relationships (Kim et al., 2020; Jang et al., 2022).
@@ 4.2 Condition Independence Based Method
Causal discovery algorithms based on conditional independence can be categorized into two distinct groups. The first category is the information-theoretic-based approach, with Optimal Causation Entropy (oCSE) (Sun et al., 2014, 2015) being a representative algorithm. oCSE is a two-step discovery algorithm explicitly designed for short time series data. The second approach is causal network-based, where the optimal causal graph is determined through statistical testing. The Peter-Clark (PC) algorithm (Kalisch and Bühlman, 2007) has gained widespread adoption and has proven effective in analyzing high-dimensional time series using causal graphs. Recognizing the potential interference of latent confounders in causal detection, corresponding approaches have been developed. A notable example is the Fast Causal Inference (FCI) algorithm (Zhang, 2008; Entner and Hoyer, 2010b; Spirtes et al., 2013), a classical method that explicitly accounts for unobserved confounders. To further enhance control over false positive rates, Runge et al. introduced the PCMCI method and its variants (Runge et al., 2019; Runge, 2020; Gerhardus and Runge, 2020) by incorporating the MCI test into the PC algorithm. PCMCI is an improvement of PC in time series, which can detect contemporaneous and time-delay effects.
@@
CDS (Fonollosa, 2019) was proposed to detect the causal direction of bivariates. This method measures the statistical characteristics of the joint distribution of marginal variance data after conditioning the bins. This algorithm has been proven to be robust as it has a high AUC in ChaLearn causal pair challenges.
@@
CD-NOD (Constraint-based causal Discovery from heterogeneous/NOnstationary Data) (Huang et al., 2020) is another framework designed to discover causal relationships in data where generating processes change over time or across domains. It detects changing local mechanisms, recovers causal structures, and estimates the driving force behind nonstationarity. This nonparametric method leverages data heterogeneity and connects nonstationarity with soft interventions, demonstrating efficacy on synthetic and real-world datasets like task-fMRI and stock market data.
@@
Here we introduce classic score-based approaches. Compared to the PC algorithm, the Greedy Equivalence Search (GES) algorithm, proposed by Chickering et al. (2002), shows enhanced robustness when dealing with nonstationary data. GRaSP (Greedy Relaxations of Sparsest Permutation) (Lam et al., 2022; Andrews et al., 2023) is designed to efficiently identify DAGs representing causal structures from observational data. It builds on permutation-based reasoning and introduces a novel operation called “tuck” to relax the assumptions required by previous methods like TSP (Triangle Sparsest Permutation) and ESP (Edge Sparsest Permutation). GRaSP consists of three tiers: GRaSP0, GRaSP1, and GRaSP2, each progressively weakening the assumptions and increasing the ability to recover sparser permutations. GRaSP2, the most relaxed form, outperforms several state-of-the-art algorithms in simulations, demonstrating scalability and accuracy for dense graphs and those with over 100 variables. To address causal discovery problems in temporal series, Pamfil et al. (2020) introduced DYNOTEARS, a novel Bayesian network learning algorithm that utilizes score constraints to ascertain the edges within the causal structure graph. DYNOTEARS is capable of effectively handling both instantaneous and delayed causality, making it a versatile tool for causal inference in time series data.
@@ 4.3 State Space Dynamic Based Method
Figure 7: Cross mapping manifold (lag = 1)
@@
4.4 Structural Equation Modelling Based Method
The three methods above are well-suited for identifying time delay causality but may not necessarily be adept at detecting instantaneous causality. On the other hand, the Structural Equation Modeling (SEM) based method represents a significant advancement and revolution in the realm of instantaneous causal discovery. This method ascertains the edges of DAG by establishing a structural equation to solve the coefficient matrix. The most basic form of the structural equation (Shimizu et al., 2006) is as follows:
@@ 4.4 Structural Equation Modelling Based Method
The Linear Non-Gaussian Acyclic Model (LiNGAM) provides a directed acyclic graph that reveals instantaneous causal relations between variables. LiNGAM assumes linear and non-Gaussian independent noise about the data generation method of the system, solving these using Independent Component Analysis (ICA) (Lee and Lee, 1998; Naik and Kumar, 2011). Since ICA algorithms typically employ FastICA and gradient-based algorithms, they may converge to local rather than global optima. To address these issues, Shimizu et al. (2011) proposed the Direct LiNGAM algorithm. Compared to ICA-LiNGAM, Direct LiNGAM produces more stable and reliable results, though it has some drawbacks. One drawback is its slower computational efficiency compared to ICA-LiNGAM. Additionally, its assumptions are relatively strict; in real-world scenarios, data generation mechanisms are often nonlinear and do not conform to its assumptions.
@@
DAGs with NO TEARS (Zheng et al., 2018) is a causal discovery method that uses continuous optimization schemes to learn the structure of Directed Acyclic Graphs (DAGs) from observational data. The name “NO TEARS” stands for “Nonlinear Optimal Transformations for Efficient and Accurate Recovery of Structure,” emphasizing its focus on tackling issues related to learning nonlinear causal relationships. The algorithm transforms the data to make it linear or Gaussian and learns a Structural Equation Model (SEM) linking features in a causal graph. It optimizes the SEM with a focus on fit and sparsity using methods like gradient descent or ADMM. Soft thresholding promotes sparsity by zeroing out edges. The process iterates until convergence criteria are met, then returns a DAG with NO TEARS, representing causal relationships.
@@ 4.5 Deep Learning Based Method
Ordering-Based Causal Discovery with Reinforcement Learning (CORL), developed by Wang et al. (2021), combines ordering-based causal discovery with reinforcement learning techniques to learn causal relationships by generating and refining an ordering of variables. The algorithm treats the problem as a sequential decision-making task, where a reinforcement learning agent arranges variables to approximate true causal relationships. A reward function provides feedback, incentivizing correct orderings and penalizing incorrect ones. The task is formulated as a Markov Decision Process (MDP), with states representing the current ordering and actions selecting the next variable. The agent is trained using reinforcement learning algorithms like Q-learning or Proximal Policy Optimization to optimize long-term rewards. Balancing exploration and exploitation is crucial. A post-processing step, such as local search, refines the ordering to improve accuracy. The algorithm ultimately returns an optimal causal graph representing the relationships between features.

### v2

@@ 2.3 Related Works
Question Data Method Evaluation
Causal Direction Pairwise Granger-Based
Condition Independent-Based Classification-Based Measures
Causal Graph I.I.D. Data State space Dynamic-Based
Structural Equation Modeling-Based
Time Series Deep Learning-Based Graph Distance-Based Measures
Hybrid Methods
Table 2: Question-Data-Method-Evaluation Template
Drawing upon the analysis mentioned above, we have amalgamated and organized the principal research directions, which are visually represented in Table 2. By employing the Question-Data-Method-Evaluation (QDME) template, the review aims to provide a clear and structured overview of the diverse areas and subtopics within the field of causal discovery research, enhancing the organization and coherence of the work.
Based on the aforementioned papers, while the algorithms and evaluations of causal discovery have become relatively comprehensive, several unresolved challenges persist, highlighting gaps in the existing body of knowledge as below.
  1. 1.
    Many articles employ outdated taxonomies and lack updates on the latest algorithms.
  2. 2.
    Most surveys emphasize theoretical analyses, often neglecting the systematic experiments necessary for quantitative assessment.
  3. 3.
    Although some articles have conducted experimental comparisons of algorithms, these studies primarily consider causal structures as their experimental factors, overlooking the characteristics of the data.
@@ 3 Survey Methodology
Figure 4 illustrates an increasing annual trend in research endeavours about causal discovery. This discernible growth can be attributed partly to the robust advancement of AI technologies in contemporary times, which has laid the groundwork for algorithmic developments facilitated by enhanced data processing capabilities.
@@ 3.1 Data Collection
“Causal Discovery” + [Survey |||| Overview |||| Review] 980 15 14
@@
Table 4: Overview of Datasets Utilised in the Reviewed Papers on Causality Research, Categorised by Causality Type, Year, Application Area, and Source.
@@
Table 5: List of Metrics Used in Reviewed Papers.
@@
Additionally, we collected several well-established packages for causal discovery algorithms, including bnlearn, pcalg, Tetrad, Causal Discovery Toolbox (CDT), CausalNex, gCastle, Tigramite, and causal-learn. bnlearn (Scutari, 2009) is an R package designed for Bayesian network learning and inference. It offers an open-source implementation of various structure learning algorithms, including constraint-based, score-based, and hybrid methods. Another R package, pcalg (Kalisch et al., 2012), integrates graphical models and causal inference techniques. It provides implementations for several widely-used causal discovery algorithms, including PC, FCI, RFCI, GES, GIES, SIMY, ARGES, and LiNGAM. Tetrad (Ramsey et al., 2018) is a Java package designed for generating and simulating data, estimating parameters, testing hypotheses, predicting outcomes, and searching causal models. CDT (Kalainathan et al., 2020) primarily focuses on discovering causal relationships from observational data, ranging from determining pairwise causal directions to full graph modeling. CausalNex (Beaumont et al., 2021) is a Python library that integrates machine learning and domain expertise for causal inference using Bayesian networks. It enables users to discover structural relationships within data, analyze complex distributions, and evaluate the effects of potential interventions. gCastle (Zhang et al., 2021b) is a causal structure learning toolchain developed by Huawei Noah’s Ark Laboratory, offering a Python library for mainstream algorithms and emerging gradient-based approaches. Tigramite (Runge et al., 2023b) is a Python package designed for time series analysis based on the PCMCI framework. It reconstructs graphical models (conditional independence graphs) from discrete or continuous time series data and generates high-quality graphical representations. causal-learn (Zheng et al., 2024b) is a Python library built upon the Java-based Tetrad causal discovery platform. The library provides modular code, enabling researchers to implement and extend their own algorithms efficiently.
3.2 Experimental Data Collection
@@ 4 Causal Discovery Algorithms
Methods, align=center [Information Theoretic-Based Approach, [oCSE (Sun et al., 2015)] ] [Causal Network-Based
Approach, align=center, s sep=2pt, l=0.9cm [Constraint-Based, [Peter-Clark (PC) (Kalisch and Bühlman, 2007)] [CD-NOD (Huang et al., 2020)] [PCMCI (Runge et al., 2019)] [Fast Causal Inference (FCI) (Entner and Hoyer, 2010)][tsFCI(Entner and Hoyer, 2010)] ] [Score-Based, [Greedy Equivalence Search (GES) (Chickering, 2002a)] [ES (Yuan and Malone, 2013)] [GRaSP (Lam et al., 2022)] [DYNOTEARS (Pamfil et al., 2020)] ] ] ] [State Space
@@
Model-Based
Methods, align=center [LiNGAM-Based, [ICA-LiNGAM (Shimizu et al., 2006)] [Direct LiNGAM (Shimizu et al., 2011)] [VARLiNGAM (Hyvärinen et al., 2010)] [RCD (Maeda and Shimizu, 2020)] [CAM-UV (Maeda and Shimizu, 2021)] ] [Additive Noise Models (ANM) (Hoyer et al., 2008), [TiMINo (Peters et al., 2013)] ] [PNL (Zhang and Hyvarinen, 2012)] [DAGs with NO TEARS (Zheng et al., 2018)] [GOLEM (Ng et al., 2020)] ] [Deep Learning-Based
Methods, align=center [Causal Generative Neural Networks (CGNN) (Goudet et al., 2018)] [DAG-GNN (Yu et al., 2019)] [Temporal Causal Discovery Framework (TCDF) (Nauta et al., 2019a)] [Amortized Causal Discovery (ACD) (Löwe et al., 2022)] [Ordering-Based Causal Discovery with Reinforcement Learning (Wang et al., 2021)] [GraNDAG (Lachapelle et al., 2019)] ] [Hybrid Methods, align=center [ARMA-LiNGAM (Kawahara et al., 2011)] [Scalable Causation Discovery Algorithm (SCDA) (Raghu et al., 2018)] [PSDR-TE (Mao and Shang, 2017)] [Information Geometric Causal Inference (IGCI) (Janzing et al., 2012)] [Non-linear SEM Estimation using Non-Stationarity (NonSENS) (Monti et al., 2020)] [Neural Granger Causality (NGC) (Tank et al., 2021b)] [NBCB and CBNB (Bystrova et al., 2024)] ] ]
Figure 5: Taxonomy for Causal Discovery (The underlined algorithms can only be applicable to Time-Series)
@@ 4.1 Granger Based Method
Our attention is directed towards several primary algorithms based on GC. Arize et al. (1993) put forth the Multivariate Granger Causality (MVGC) analysis method to overcome the limitations of Pairwise Granger Causality (PWGC), which can only deal with bivariate data. However, ensuring linearity in real-world datasets can present a significant challenge. To address this issue, Chen et al. (2004) introduced an approach known as Extended Granger Causality (EGC), specifically designed to handle nonlinear data. An alternative model catering to nonlinear time series is the Kernel Granger Causality (KGC) method (Marinazzo et al., 2008; Liao et al., 2009; Marinazzo et al., 2011), showcasing notable attributes such as high accuracy and flexibility. When confronted with continuous time series data, Hu et al. (2014) proposed the Copula Granger method, which is capable of uncovering nonlinear and higher-order causal relationships (Kim et al., 2020; Jang et al., 2022).
Although Granger-based causality method has a long history of development, it still struggles to handle complex causal relationships. Its main drawback is the inability to identify latent confounders and instantaneous causal effects. Therefore, the Granger-based method is often combined with other methods to achieve mutual development.
@@ 4.2 Condition Independence Based Method
Causal discovery algorithms based on conditional independence can be categorized into two distinct groups. The first category is the information-theoretic-based approach, with Optimal Causation Entropy (oCSE) (Sun et al., 2014, 2015) being a representative algorithm. oCSE is a two-step discovery algorithm explicitly designed for short time series data. The second approach is causal network-based, where the optimal causal graph is determined through statistical testing. The Peter-Clark (PC) algorithm (Kalisch and Bühlman, 2007) has gained widespread adoption and has proven effective in analyzing high-dimensional time series using causal graphs. Recognizing the potential interference of latent confounders in causal detection, corresponding approaches have been developed. A notable example is the Fast Causal Inference (FCI) algorithm (Zhang, 2008; Entner and Hoyer, 2010; Spirtes et al., 2013), a classical method that explicitly accounts for unobserved confounders. To further enhance control over false positive rates, Runge et al. introduced the PCMCI method and its variants (Runge et al., 2019; Runge, 2020; Gerhardus and Runge, 2020) by incorporating the MCI test into the PC algorithm. PCMCI is an improvement of PC in time series, which can detect contemporaneous and time-delay effects.
@@
Conditional Distribution Similarity Statistic (CDS) (Fonollosa, 2019) was proposed to detect the causal direction of bivariates. This method measures the statistical characteristics of the joint distribution of marginal variance data after conditioning the bins. This algorithm has been proven to be robust as it has a high AUC in ChaLearn causal pair challenges.
@@
Constraint-based causal Discovery from heterogeneous/NOnstationary Data (CD-NOD) (Huang et al., 2020) is another framework designed to discover causal relationships in data where generating processes change over time or across domains. It detects changing local mechanisms, recovers causal structures, and estimates the driving force behind nonstationarity. This nonparametric method leverages data heterogeneity and connects nonstationarity with soft interventions, demonstrating efficacy on synthetic and real-world datasets like task-fMRI and stock market data.
@@
Here we introduce classic score-based approaches. Compared to the PC algorithm, the Greedy Equivalence Search (GES) algorithm, proposed by Chickering et al. (2002a), shows enhanced robustness when dealing with nonstationary data. Greedy Relaxations of Sparsest Permutation (GRaSP) (Lam et al., 2022; Andrews et al., 2023) is designed to efficiently identify DAGs representing causal structures from observational data. It builds on permutation-based reasoning and introduces a novel operation called “tuck" to relax the assumptions required by previous methods like Triangle Sparsest Permutation (TSP) and Edge Sparsest Permutation (ESP). GRaSP consists of three tiers: GRaSP0, GRaSP1, and GRaSP2, each progressively weakening the assumptions and increasing the ability to recover sparser permutations. GRaSP2, the most relaxed form, outperforms several state-of-the-art algorithms in simulations, demonstrating scalability and accuracy for dense graphs and those with over 100 variables. To address causal discovery problems in temporal series, Pamfil et al. (2020) introduced DYNOTEARS, a novel Bayesian network learning algorithm that utilizes score constraints to ascertain the edges within the causal structure graph. DYNOTEARS is capable of effectively handling both instantaneous and delayed causality, making it a versatile tool for causal inference in time series data.
Compared to Granger-based methods, conditional independence-based methods can handle more complex data scenarios, such as high-dimensional data, instantaneous causality, and latent variables. However, these methods generally require the faithfulness assumption and have limitations in determining causal direction, i.e., some causal links remain unoriented. Despite these drawbacks, they are well-suited for identifying causal skeleton graphs.
@@ 4.3 State Space Dynamic Based Method
Figure 6: Cross mapping manifold (lag = 1) (Javier, 2021)
@@
The most significant advantage of state space dynamic based methods is their efficacy in deterministic systems, making them the preferred choice for such specialized data scenarios. However, their primary disadvantage lies in their limited applicability and difficulty in handling datasets with time-variant noises.
4.4 Structural Equation Model Based Method
The three methods above are well-suited for identifying time delay causality but may not necessarily be adept at detecting instantaneous causality. On the other hand, the Structural Equation Model (SEM) based method represents a significant advancement and revolution in the realm of instantaneous causal discovery. This method ascertains the edges of DAG by establishing a structural equation to solve the coefficient matrix. The most basic form of the structural equation (Shimizu et al., 2006) is as follows:
@@ 4.4 Structural Equation Model Based Method
The Linear Non-Gaussian Acyclic Model (LiNGAM) provides a directed acyclic graph that reveals instantaneous causal relations between variables. LiNGAM assumes linear and non-Gaussian independent noise about the data generation method of the system, solving these using Independent Component Analysis (ICA) (Lee and Lee, 1998; Naik and Kumar, 2011). Since ICA algorithms typically employ FastICA and gradient-based algorithms, they may converge to local rather than global optima. To address these issues, Shimizu et al. (2011) proposed the Direct LiNGAM algorithm. Compared to ICA-LiNGAM, Direct LiNGAM produces more stable and reliable results, though it has some drawbacks. One drawback is its slower computational efficiency compared to ICA-LiNGAM. Additionally, its assumptions are relatively strict; in real-world scenarios, data generation mechanisms are often nonlinear and do not conform to its assumptions. Hyvärinen et al. (2013) proposed a measure for determining pairwise causal direction based on likelihood ratio tests. The objective of the study was to develop a method based on SEM that performs more effectively on real-world data with limited sample sizes. In simulations using brain imaging (fMRI) data, this method demonstrated significantly better performance compared to ICA-LiNGAM and other approaches.
@@
DAGs with NO TEARS (Zheng et al., 2018) is a causal discovery method that uses continuous optimization schemes to learn the structure of Directed Acyclic Graphs (DAGs) from observational data. The name “NO TEARS” stands for “Nonlinear Optimal Transformations for Efficient and Accurate Recovery of Structure,” emphasizing its focus on tackling issues related to learning nonlinear causal relationships. The algorithm transforms the data to make it linear or Gaussian and learns a Structural Equation Model (SEM) linking features in a causal graph. It optimizes the SEM with a focus on fit and sparsity using methods like gradient descent or ADMM. Soft thresholding promotes sparsity by zeroing out edges. The process iterates until convergence criteria are met, then returns a DAG with NO TEARS, representing causal relationships. However, Kaiser and Sipos (2021) analyzed the lack of scale invariance in the NOTEARS algorithm and concluded that this limitation makes NOTEARS unsuitable for identifying true causal relationships from data.
Regression Error based Causal Inference (RECI) (Blöbaum et al., 2018) addresses the problem of inferring the causal relationship between two variables by comparing the least-squares errors of predictions in both possible causal directions. Blöbaum emphasize that RECI can have a significantly lower computational cost than ANM, while delivering comparable or even superior results. Additionally, RECI is straightforward to implement and apply.
@@
The advantage of Structural Equation Model based methods is their applicability to a wide range of data types. LiNGAM and its variant algorithms can handle i.i.d. data, time series, instantaneous causality, hidden confounders, mixed data, and other data types without requiring faithfulness assumptions. However, these methods predominantly rely on linear relationships, and only a few algorithms are capable of handling nonlinear relations.
@@ 4.5 Deep Learning Based Method
Ordering-Based Causal Discovery with Reinforcement Learning (CORL), developed by Wang et al. (2021), combines ordering-based causal discovery with reinforcement learning techniques (Zhu et al., 2019) to learn causal relationships by generating and refining an ordering of variables. The algorithm treats the problem as a sequential decision-making task, where a reinforcement learning agent arranges variables to approximate true causal relationships. A reward function provides feedback, incentivizing correct orderings and penalizing incorrect ones. The task is formulated as a Markov Decision Process (MDP), with states representing the current ordering and actions selecting the next variable. The agent is trained using reinforcement learning algorithms like Q-learning or Proximal Policy Optimization to optimize long-term rewards. Balancing exploration and exploitation is crucial. A post-processing step, such as local search, refines the ordering to improve accuracy. The algorithm ultimately returns an optimal causal graph representing the relationships between features.
@@
The advantage of deep learning based methods lies in their capacity to handle datasets with large sample sizes and numerous variables. However, their drawbacks include long running times, low efficiency, and suboptimal performance on short time series.